A 211 g ball is tied to a string. It is pulled to an angle of 6.4 ∘ and released to swing as a pendulum. A student with a stopwatch finds that 15 oscillations take 15 s .

How long is the string?
Answer #1

m=0.211 kg

T= 15/15= 1 s

g= 9.8 m/s^2

Formula for time period of a pendulum

T = 2*3.14*sqrt(L/g)

L=g*(T/6.28)^2=0.2485 meters= 24.85 cm